OpenStudy (anonymous):

Someone help please. An expression is shown below: f(x) = -16x2 + 24x + 16 Part A: What are the x-intercepts of the graph of f(x)? Show your work. Part B: Is the vertex of the graph of f(x) going to be a maximum or minimum? What are the coordinates of the vertex? Justify your answers and show your work. Part C: What are the steps you would use to graph f(x)? Justify that you can use the answers obtained in Part A and Part B to draw the graph.

OpenStudy (anonymous):

@triciaal @zepdrix @mathmale @mathstudent55

OpenStudy (anonymous):

@Directrix @Hero @ParthKohli @pooja195 @Mehek14 @mayankdevnani @Jaynator495

zepdrix (zepdrix):

Sup :d

OpenStudy (anonymous):

Sup

OpenStudy (anonymous):

sup sup sup sup sup sup sup

zepdrix (zepdrix):

The function crosses the x-axis when.... the y-coordinate is zero. Hopefully that makes sense. So to find our x-intercepts, where the function crosses the x-axis, we'll replace our y with zero. And then solve for x.

zepdrix (zepdrix):

\[\large\rm f(x)=-16x^2+24x+16\]\[\large\rm ~~~~0=-16x^2+24x+16\]

zepdrix (zepdrix):

The next few steps should feel familiar, it's what you've been doing a lot of, factoring :)

OpenStudy (anonymous):

@zepdrix I gotta go, be back in 5 minutes. sorry lol

OpenStudy (anonymous):

@zepdrix Dude, really sorry, but I have to leave for like 6 hours Sorry.

zepdrix (zepdrix):

k c:

OpenStudy (anonymous):

@zepdrix I'm back lol

OpenStudy (anonymous):

@zepdrix So I need your help now. I get that y has to be 0 and then solve for x, but what then?

OpenStudy (anonymous):

@zepdrix All right, ill be back in 15 - 20 minutes

zepdrix (zepdrix):

then what? then do it, solve for x :o get your factors.

OpenStudy (anonymous):

Alright, I'll try

OpenStudy (mathmale):

@needhelpstudying: Please, would you plan your time so that you have enough time in which to complete a whole problem before excusing yourself? Constantly excusing yourself could be interpreted to be disrespect for the person helping you. zepdrix is correct in asking you to solve the equation \[-16x^2+24x+16 = 0.\]

OpenStudy (mathmale):

What is the largest number (greatest common factor) that will divide into 16 and 24 with no remainder? Factor out the GCF. Rewrite the equation in this factored form. Which methods for solving a quadratic equation do you know: For starters: completing the square, factoring, graphing, quadratic formula.

OpenStudy (anonymous):

@mathmale Sorry, some family friends came to my house, my parents didn't really give me any warning at all lol. Should I be factoring this by grouping? −16x2+24x+16

OpenStudy (anonymous):

@zepdrix If ur online, I can start now

Nnesha (nnesha):

as mentioned above first always look for GCF ( greatest common factor ) is there any thing common ? if yes take it out first and then factor it

zepdrix (zepdrix):

eyy broski. grouping? naw not just yet. do that ^

zepdrix (zepdrix):

They're all even numbers, so they at least have a 2 in them. Anything larger as a factor in each?

OpenStudy (anonymous):

@zepdrix Sorry, laptop shut off. Alright. GCF Of 8 and 24 is 8. so GCF would be 8x. Factor it out: -16x2 + 24x + 16 8x(-2x + 3) + 16 Correct?

zepdrix (zepdrix):

No no no, we're looking to factor something out of EVERY TERM to start this one out.

zepdrix (zepdrix):

The last term, 16, has no x, so we can't pull 8x out. We'll have to result to pulling just 8 out of everything, ya?

zepdrix (zepdrix):

resolve*

zepdrix (zepdrix):

\[\large\rm -16x^2+24x+16\]\[\large\rm 8(-2x^2+3x+2)\]Understand that step?

OpenStudy (anonymous):

@zepdrix Yeah, I get it now. I kinda ignored that last term, sorry

OpenStudy (anonymous):

Yeah, so i get that step

zepdrix (zepdrix):

These usually work out a little nicer when the squared term is POSITIVE. So let's apply another step, hope this one doesn't confuse you too much. We're going to factor -1 out of each term. Essentially what that does is, it just changes the `sign` of everything.\[\large\rm -8(2x^2-3x-2)\]

OpenStudy (anonymous):

HOORAY FOR CHANGING SIGNS!

OpenStudy (anonymous):

Wait, so what do you do to the -8 that's already outside the parantheses?

zepdrix (zepdrix):

Well, recall that we were letting this entire thing equal 0.\[\large\rm 0=-8(2x^2-3x-2)\]So we can do something clever here. We can divide both sides by -8, our equation becomes this:\[\large\rm 0=2x^2-3x-2\]

OpenStudy (anonymous):

NEVER NEVER NEVER NEVER NEVER BTW It's -2x2

OpenStudy (anonymous):

@zepdrix Hello?????????????????????

zepdrix (zepdrix):

whut?

zepdrix (zepdrix):

The -2x^2 became 2x^2 when we took the negative out of everything.

OpenStudy (anonymous):

Wait You wrote 0=2x2−3x−2 Should it be 0 / 8=2x2−3x−2

zepdrix (zepdrix):

Yes :) And what is 0 divided by -8?

OpenStudy (anonymous):

@Needhelpstudying is amazing?

OpenStudy (anonymous):

lololololololol 0 - 8 is ZEPDRIX!

Nnesha (nnesha):

take it serious.

OpenStudy (anonymous):

0 / -8 = 0. Then the shark came and bit off my head. The End.

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@ganeshie8 Can you help?

ganeshie8 (ganeshie8):

which part ?

OpenStudy (anonymous):

Pretty much everything

ganeshie8 (ganeshie8):

\[\large\rm 0=2x^2-3x-2\] have you tried to factor this ?

OpenStudy (anonymous):

I don't think anything can be factored from it.

ganeshie8 (ganeshie8):

pick two numbers such that 1) product is "-4" 2) sum is "-3"

OpenStudy (anonymous):

GAAH How do you do those problems again?

OpenStudy (anonymous):

@ganeshie8 Is there any formula for finding the answer, or is it just guess and check?

OpenStudy (anonymous):

Factors of -4 -4, 1 -1, 4 2, -2

OpenStudy (anonymous):

The answer is -4, 1. Correct @ganeshie8 ?

ganeshie8 (ganeshie8):

Yes, now rewrite the middle term -3x as -4x + 1x

OpenStudy (anonymous):

\[0=2x^2−x−2\]

ganeshie8 (ganeshie8):

\[\large\rm 0=2x^2\color{red}{-3x}-2\] is samea as \[\large\rm 0=2x^2\color{red}{-4x+1x}-2\]

ganeshie8 (ganeshie8):

Next, group first two terms group last two terms

OpenStudy (anonymous):

Thats strange, when I wrote the equation it didn't include everything I typed....

OpenStudy (anonymous):

0=(2x2−4x) + (1x−2)

ganeshie8 (ganeshie8):

happens... so what is the common factor in first two terms ?

OpenStudy (anonymous):

2x?

OpenStudy (anonymous):

and 1 for the second term?

ganeshie8 (ganeshie8):

Yes, pull that out

OpenStudy (anonymous):

alright 0=2x(x − 2) + 1(x − 2)

OpenStudy (anonymous):

0 = (2x + 1)(x − 2)

OpenStudy (anonymous):

Correct?

ganeshie8 (ganeshie8):

Looks good. For that product to be really equal to 0, we must necessarily have either one of the factors equal to 0 : 0 = 2x + 1 or 0 = x - 2

ganeshie8 (ganeshie8):

you can solve x

OpenStudy (anonymous):

x = -1/2 x = 2 ?

ganeshie8 (ganeshie8):

Yep! those are the x intercepts. The graph of given function f(x) cuts the x axis at -1/2 and 2

OpenStudy (anonymous):

WOOHOOO!!!

OpenStudy (anonymous):

I just realized after all this i've only solved part a. :(

ganeshie8 (ganeshie8):

Lets graph f(x) quick and verify part a

OpenStudy (anonymous):

K, desmos?

OpenStudy (anonymous):

Yeah, it's correct. I just checked.

ganeshie8 (ganeshie8):

good

OpenStudy (anonymous):

On to part b

ganeshie8 (ganeshie8):

For part b : The leading coefficient of f(x) is -16 which is negative. Therefore the vertex of f(x) is going to be a minimum.

OpenStudy (anonymous):

wait wait wait, what is a vertex?

ganeshie8 (ganeshie8):

Look at the graph again. How does it look ? Do you see any "hill top" ?

ganeshie8 (ganeshie8):

In our case, the point on the top of the hill is called the vertex point.